OpenStudy (studyhoney):

can someone help me learn equations? (for algebra 1.. I have to take a state test so I want to prepare)

OpenStudy (amorfide):

if you can show me an example of the types of equations you need to learn then sure

OpenStudy (amorfide):

I say this because your algebra 1 may be different to mine

OpenStudy (studyhoney):

okay here is an example: -4(x + 10) - 6 = -3(x - 2)

OpenStudy (amorfide):

ooo! okay

OpenStudy (amorfide):

suppose you have 4+5=9

OpenStudy (amorfide):

if I were to multiply both sides by 6 I would end up with 6(4+5)=6(9) right?

OpenStudy (studyhoney):

this isnt the equation i posted right?

OpenStudy (amorfide):

no it is not

OpenStudy (amorfide):

I am starting from the basics, and we will work our way u p

OpenStudy (studyhoney):

okay yeah i got that

OpenStudy (studyhoney):

you just added the 6 right?

OpenStudy (amorfide):

okay so when we do 6(4+5)=6(9) this is \[(6 \times 4) + (6 \times 5)=6 \times 9\]

OpenStudy (studyhoney):

okay...

OpenStudy (amorfide):

so let's show this in a more general form, using algebra

OpenStudy (amorfide):

suppose we have a+b=c then we multiply it by a number, D we get D(a+b)=D(c) Da + Db = Dc so if I give you -4(a+6)=-3 can you attempt to solve for a

OpenStudy (studyhoney):

im not exactly sure how to solve that

OpenStudy (amorfide):

-4(a+6) to multiply this out of the bracket  multiply the first term what do you get?